OpenStudy (anonymous):

it is known that sin2θ=2sinθcosθ prove show that sin4θ=sinθcosθcos2θ

OpenStudy (anonymous):

you want \[sin(4\theta)=4sin(\theta)cos(\theta)cos(2\theta)\]

OpenStudy (raden):

sin4θ=sinθcosθcos2θ ??? looks this is not an identity

OpenStudy (raden):

have u typed correctly ?

OpenStudy (anonymous):

sorry there's a typo it should be sin(4θ)=4sin(θ)cos(θ)cos(2θ)

OpenStudy (anonymous):

That is \[sin(4\theta)=2sin(2\theta)cos(2\theta)=4sin(\theta)cos(\theta)cos(2\theta)\]

OpenStudy (raden):

sin2θ=2sinθcosθ (known) sin4θ = sin2(2θ) = 2sin2θcos2θ sin4θ = 2(2sinθcosθ)cos2θ sin4θ = 4sinθcosθcos2θ

OpenStudy (anonymous):

thank you!